# Language-Guided World Models [Uncaptioned image] A Model-Based Approach to AI Control

### v1

@@ 4 The Messenger-Wm benchmark: Evaluating the compositional generalizability of world models
Our benchmark is built on top of the Messenger environment and dataset (Hanjie et al., 2021). Messenger is a role-playing game that takes place in a $10\times 10$ grid world. A player interacts with entities having one of three roles: message, goal, or enemy. The objective of the player is to acquire the message and deliver it to the goal while avoiding the enemy. In addition to the role, each entity is assigned an identity among twelve possibilities (mage, airplane, orb, etc.) and a movement pattern (chasing the agent, fleeing from the agent, immobile).
Each game features a combination $\{(i^{c},v_{\textrm{move}}^{c},v_{\textrm{role}}^{c})\}_{c=1}^{3}$ where $i^{c}$ denotes the identity of the entity $c$, $v_{\textrm{move}}^{c}$ its movement pattern, and $v_{\textrm{role}}^{c}$ its role. There is exactly one entity for each role. The movement patterns and roles of the entities are the latent factors that dictate the environment dynamics; the identity is irrelevant, but observable.
@@
A game’s manual consists of three descriptions. The authors of Messenger collected a data set of 5,316 language descriptions, each of which describes a combination of identity, role, and movement. Manual descriptions employ various linguistic expressions for each identity, role, or movement pattern (e.g., an airplane can be mentioned as a “plane”, “jet”, or “airliner”), making it nontrivial to parse the referred identities and attributes. Evaluation descriptions are completely unseen during training.
Evaluating compositional generalization.
We create three levels of compositional generalization:
  • NewCombo (easy). Each game features a combination of three identities that were never seen together in a training game. However, the attributes (role and movement pattern) of each identity are the same as during training.
@@
To generate trajectories, we implement rule-based behavior policies that execute various intentions: act randomly, avoid the enemy, suicide (go to the enemy), obtain the message, and win the game (obtain the message and deliver it to the goal). We generate a total of 100K trajectories for training, each of which is generated by rolling out a uniformly randomly chosen rule-based policy. More details of the data are given in Appendix A.
Challenging world models.
Modeling the dynamics of a new environment without interactions is daunting, especially when the training data deliberately present strong spurious correlations to deceive models. When a world model is deployed for planning or learning, it must make decisions based on self-generated histories. Every error made in one time step propagates to all future time steps. Our hardest evaluation setting ensures that it is highly unlikely that a model can accurately predict the first few time steps without understanding the manual. In general, our evaluation is more comprehensive than the stage-two evaluation of the original Messenger paper, which does not construct different levels of compositional generalization, and is more difficult than that of Lin et al. (2023), which does not test for generalization.
@@ 5.1 Baselines
We compare our proposed model, which we call EMMA-LWM, with the following models:
@@
    Observational world model learns from only observations. It is identical to EMMA-LWM except that the manual representation $\bm{m}^{\textrm{enc}}$ is a zero matrix rather than a representation of the text manual;
@@
    Standard is the encoder-decoder Transformer model introduced by Vaswani et al. (2017) with multi-headed cross-attention between the decoder and the encoder. Similarly to EMMA-LWM, the model uses BERT to initially encode the manual into hidden vectors. The encoder applies self-attention to the hidden vectors of each description separately, instead of joining all vectors into a sequence and applying self-attention on it;
@@
    GPTHard implements the ChatGPT-based hard attention approach described in § 3.2;
@@
    OracleParse assumes oracle semantic parsing. It is similar to the GPTHard approach except that the manual is parsed into canonical identity and attribute names. For example, a description “the crucial target is held by the wizard and the wizard is fleeing from you” is converted into “mage fleeing goal” for the model.
We train all models using the AdamW optimizer (Loshchilov & Hutter, 2017) for $10^{5}$ iterations with a learning rate of $10^{-4}$ and a batch size of 32. For further details, please refer to Appendix C.
Table 1: Cross entropy losses ($\downarrow$) of different models on test ground-truth trajectories. Note that the minimal losses are non-zero because Messenger is stochastic. We run each model with five different random seeds, selecting the final checkpoint for each seed based on the loss in the development NewAll split. We report the mean losses with 95% t-value confidence intervals. The bold number in each column indicates the best non-oracle mean. EMMA-LWM demonstrates robustness in the more difficult splits (NewAttr and NewAll), approaching the performance of an oracle model with a perfect semantic-parsing capability.
@@ 5.2 Intrinsic evaluation
Table 1 shows the cross-entropy losses of all models on ground-truth trajectories sampled from the true environment dynamics. In the more difficult NewAttr and NewAll splits, our EMMA-LWM model consistently outperforms all baselines, nearing the performance of the OracleParse model. As expected, the Observational model is easily fooled by spurious correlations between identity and attributes, and among attributes. A specific example is illustrated in Figure 3. There, the Observational model incorrectly captures the movement of the whale and the queen. It also mistakenly portrays the whale as an enemy, whereas, in fact, the entity holds the message. In contrast, EMMA-LWM is capable of interpreting the previously unseen manual and accurately simulates the dynamics of the environment.
The performance of the Standard model is sensitive to initialization; in some runs, it performs as well as EMMA-LWM, but in others it performs as badly as Observational. A plausible explanation is that the model’s attention mechanism lacks sufficiently strong inductive biases to consistently find generalizable solutions. Our results agree with previous work on the lack of compositional generalizability of Transformers, which is often remedied by adding numerous forms of inductive bias (Keysers et al., 2020; Jiang & Bansal, 2021; Chaabouni et al., 2021; Dziri et al., 2023).
Another interesting finding is that the GPTHard model does not perform as well as we expected. As a reminder, this model relies on ChatGPT to parse identities from descriptions and only needs to learn to extract attributes. Its underperformance compared to EMMA-LWM can be attributed to (i) the imperfection of ChatGPT in identifying identities in descriptions (its accuracy is around 90%; see Appendix A) and (ii) the fact that EMMA-LWM jointly learns to extract both identity and attribute words, which may be more effective than learning to extract only attribute words.
@@
In this evaluation, for each world model and test trajectory, we reset the model to the initial state of the trajectory and sequentially feed actions in the trajectory to the model until it predicts the end of the episode. This process generates an imaginary trajectory. We refer to the evaluation trajectory as the real trajectory. We compute precisions of predicting non-zero rewards ($r\neq 0$) and terminations ($d=1$). To evaluate movement prediction, we compare the distances from the player to an entity in the real and imaginary trajectories. Concretely, let $\delta_{c,t}^{\text{real}}$ and $\delta_{c,t}^{\text{imag}}$ be the Hamming distances from the player to entity $c$ at the $t$-th time step in a real trajectory $\tau_{\textrm{real}}$ and an imaginary trajectory $\tau_{\textrm{imag}}$, respectively. We calculate the average difference in a specific time step: $\Delta_{\text{dist}}=\frac{1}{|\mathcal{D}_{\textrm{eval}}|}\sum_{\tau_{\textrm{real}}\in\mathcal{D}_{\textrm{eval}}}\frac{1}{T_{\textrm{min}}}\sum_{t=1}^{T_{\textrm{min}}}|\delta_{c,t}^{\text{real}}-\delta_{c,t}^{\text{imag}}|$ where $\mathcal{D}_{\textrm{eval}}$ is an evaluation split, $T_{\textrm{min}}=\min(|\tau_{\textrm{real}}|,|\tau_{\textrm{imag}}|)$, and $\tau_{\textrm{imag}}$ is generated from $\tau_{\textrm{real}}$ For example, for a chasing entity, $\delta_{c,t}^{\text{real}}$ decreases as $t$ increases. If a model mistakenly predicts the entity to be immobile, $\delta_{c,t}^{\text{imag}}$ remains a constant as $t$ progresses. In this case, $\Delta_{\text{dist}}$ is non-negligible, indicating an error. All evaluation metrics are given in Table 2. The ordering of the models is similar to that in the evaluation with ground-truth trajectories. EMMA-LWM is still superior to all baselines in all metrics.
@@ 5.3 Application: agents that discuss plans with humans
Table 3: Average returns ($\uparrow$) in real environments of policies trained with imaginary imitation learning using world models. For each world model type, we randomly choose one out of the five checkpoints mentioned in Table 1. Experiments are conducted in 30 environments for each test split. For each environment and learned policy, we compute the average return over 48 runs. We report the means of the average returns in the 30 environments with 95% t-value confidence intervals. Bold numbers indicate the best non-oracle means in the corresponding settings. Results of all models are available in Appendix F.
@@
In this section, we showcase the effectiveness of our LWM in facilitating plan discussion between an agent and its human supervisor, which consequently improves the interpretability, safety, and performance of the agent.
We imagine an agent ordered to perform a task in a previously unseen environment (0(a)). Letting the agent perform the task immediately would be extremely risky because of its imperfect knowledge of the environment. Implementing a world model enables the agent to imagine a solution trajectory and present it to a human as a plan for review. Conveying plans as trajectories helps the human envision the future behavior of the agent in the real world. Furthermore, the human can improve this behavior by providing feedback for the agent to revise its plan.
The human can provide feedback that rectifies the agent’s actions. However, enhancing the agent’s plan requires correcting not only its actions, but also its world model. Although the agent can update its world model on its own by collecting extra experiences in the real environment, doing so would defy the very purpose of having a plan discussion, which is to prevent the agent from acting recklessly in the real world. An LWM resolves this problem by enabling the world model to learn from language feedback given by the human. An observational world model would require much more effort from the human to communicate feedback; they have to be able to generate observations in the same format as those in the agent’s plans (e.g., they have to draw grids). Furthermore, abstract concepts (e.g., “slippery”) may not be precisely conveyed through a few images.
Simulating this scenario, we implement a randomly initialized policy (the agent) learning to solve a Messenger environment with an expert policy (the human supervisor). The learning policy is forbidden to interact with the real environment until test time. Learning is entirely imaginary: all trajectories are generated using a world model. Importantly, the world model was never trained on any data collected in the real environment, simulating the fact that the environment is previously unseen. After training, we run the learned policy in the real environment and record its average return. We consider two types of action-correcting feedback: in the online imitation learning setting (Ross et al., 2011), the expert corrects actions in agent-generated trajectories; in the filtered behavior cloning setting, the expert presents demonstrative trajectories. In the latter setting, before each learning episode, we add a set of demonstrative trajectories to a buffer and retain only those that achieve the highest returns in the buffer for the policy to imitate. We experiment with a near-optimal expert and a suboptimal expert.
We present the results in Table 3. Learning with the Observational world model amounts to the case where the human provides only action-correcting feedback and cannot adapt the world model. Meanwhile, learning with EMMA-LWM represents the case where the human can use language feedback to improve the world model. In all evaluation settings, we observe significant improvements in the average return of policies that adopt our EMMA-LWM. There are still considerable gaps compared to using the OracleParse model, indicating that our model still has room for improvement.

### v2

@@ 3.3 The Messenger-WM benchmark
The game takes place in a $10\times 10$ grid world. A player interacts with entities of three roles: message, goal, and enemy. We use the stage-two version of the game, in which there are three entities, one of each role, in a game instance. In addition to the role, each entity is assigned an identity among twelve possibilities (mage, airplane, orb, etc.) and a movement pattern (chasing the agent, fleeing from the agent, immobile). The objective of the player is to acquire the message and deliver it to the goal while avoiding the enemy. Fetching the message is awarded 0.5 points and delivering it to the goal adds another point. If the player collides with the enemy or reaches the goal without carrying the message, the game ends, and the player receives -1 points.
@@
A game’s manual consists of three descriptions corresponding to the three entities. Messenger provides a dataset of 5,316 language descriptions, each of which describes a combination of identity, role, and movement. The descriptions employ various linguistic expressions for each identity, role, or movement pattern (e.g., an airplane can be mentioned as a “plane”, “jet”, or “airliner”), making it non-trivial to interpret.
Evaluation settings.
To test for compositional generalization, we construct three evaluation settings, ordered in increasing degree of difficulty:
  • NewCombo (easy). Each game features a combination of three identities that were never seen together in a training game. However, the role and movement pattern of each identity are the same as during training.
@@
To generate trajectories, we implement rule-based behavior policies that execute various intentions: act randomly, avoid the enemy, suicide (go to the enemy), obtain the message, and win the game (obtain the message and deliver it to the goal). We generate a total of 100K trajectories for training, each of which is generated by rolling out a uniformly randomly chosen rule-based policy. More details of the data are given in Appendix B. Our evaluation is more comprehensive than the original Messenger paper’s evaluation, which does not construct different levels of compositional generalization, and is more difficult than the setting of Lin et al. (2024), which does not concern generalization.
To succeed in Messenger-WM, a model must be able to understand the non-trivial concepts represented by the attributes. For example, the concept of “chasing” involves planning actions to reduce the distance between two entities. The model must also capture the independence of the attributes, despite observing correlations in the training data (e.g., the “mage” is never immobile during training). Finally, to reflect the locality of the attributes, the model needs to learn a representation that disentangles the entities and to route attributes to the right entities. For example, the movement of one entity should not influence that of another. These are among the difficult, under-explored problems in machine learning, making Messenger-WM a respectable research challenge. We will empirically show that the state-of-the-art Transformer architecture struggles to perform well on the benchmark, suggesting that it may be insufficient for tackling more complex world-modeling problems.
4 Modeling approach
State representation.
In Messenger, a state $s$ is represented by an $H\times W$ grid with $C$ channels (an $H\times W\times C$ tensor), where each channel corresponds to an entity. In each channel $c$, there is a single non-zero cell $s(h,w,c)$ that represents the identity of the entity. The position of this cell is the location of the entity in the grid. We note that this is an idealized representation that disentangles the entities. Even so, the problem remains challenging, as the model needs to recognize attributes mentioned in the manual and associate them with the right entity token. This requires a special attention mechanism, which we will introduce shortly. Meanwhile, learning entity-disentangled representations for pixel-based environments remain an open problem, which we defer to future work.
World modeling as sequence generation.
Our model (illustrated in 1(b)) is an encoder-decoder Transformer (Vaswani et al., 2017) which encodes a manual $\bm{\ell}$ and decodes a trajectory $\tau$. We transform the trajectory into a long sequence of tokens and train the model as a sequence generator.
Concretely, our model processes a data point $(\tau,\bm{\ell})$ as follows. For the manual =$\bm{\ell}=\{l_{i}\}_{i=1}^{N}$, we first use a pre-trained BERT model to convert each description $l_{i}$ into a sequence of hidden vectors. We feed each sequence to a Transformer encoder, which outputs a tensor $\bm{m}^{\textrm{enc}}$ of size $N\times L\times D$, where $N=C$ is the number of descriptions, $L$ is the maximum number of words in a description, and $D$ is the hidden size.
For the trajectory, we convert each tuple $(a_{t-1},s_{t},r_{t},d_{t})$ into a token block $B_{t}$. The first action $a_{0}$ is set to be a special <s> token. Each state $s_{t}$ is mapped to $3C$ tokens $(i_{t}^{1},h^{1}_{t},w^{1}_{t},\cdots,i^{C}_{t},h^{C}_{t},w^{C}_{t})$, which represents each of the $C$ entities by its identity $i$ followed by its location $(h,w)$. The real-valued reward $r_{t}$ is discretized into an integer label, and the termination signal $d_{t}$ is translated into a binary label. In the end, $B_{t}$ consists of $33C+$ 3 tokens $(a_{t-1},i_{t}^{1},h^{1}_{t},w^{1}_{t},\cdots,i^{C}_{t},h^{C}_{t},w^{C}_{t},r_{t},d_{t})$. Finally, we concatenate all $T$ blocks in the trajectory into a sequence of $T\times(3C+3)$ tokens, embed them into a $T\times(3C+3)\times D$ tensor, and add positional embeddings. We will use bold notation (e.g., $\bm{a}$, $\bm{i}$) to refer to the resultant embeddings of the tokens.
Entity mapper with multi-modal attention.
We implement a variant of EMMA (Hanjie et al. (2021)) that first identifies the description that mentions each entity and extracts from it words corresponding to the attributes of the entity. From the tensor $\bm{m}_{n}^{\textrm{enc}}$ computed by the encoder, we generate a key tensor $\bm{m}^{\textrm{key}}$ and a value tensor $\bm{m}^{\textrm{val}}$, both of which are of size $N\times L\times D$, where
$\bm{m}^{\textrm{key}}_{n}$ $=\texttt{Softmax}(\texttt{Linear}_{\textrm{key}}(\bm{m}_{n}^{\textrm{enc}})^{\top})\bm{m}_{n}^{\textrm{enc}}$
$\bm{m}^{\textrm{val}}_{n}$ $=\texttt{Softmax}(\texttt{Linear}_{\textrm{val}}(\bm{m}_{n}^{\textrm{enc}})^{\top})\bm{m}_{n}^{\textrm{enc}}$ (1)
for $1\leq n\leq N$. Here, $\texttt{Linear}^{D\rightarrow 1}_{\textrm{key}}$ and $\texttt{Linear}^{D\rightarrow 1}_{\textrm{val}}$ are linear layers that transform the input’s last dimension from $D$ to 1, and $\texttt{Softmax}(\cdot)$ applies the softmax function to the last dimension. Intuitively, we want each $\bm{m}^{\textrm{key}}_{n}$ to retain words that signal the identity of the entity mentioned in the $n$-th description (e.g., ferry, plane, researcher), and $\bm{m}^{\textrm{val}}_{n}$ to retrieve words depicting the other attributes (e.g., approaching, deadly, fleeing).
Let $\bm{i}^{c}_{t}$ be the embedding of the identity of entity $c$. We perform a dot-product attention with $\bm{i}^{c}_{t}$ as the query, $\bm{m}^{\textrm{key}}$ as the set of keys, and $\bm{m}^{\textrm{val}}$ as the set of values to compute the attribute features of $c$
The features are added to the identity tokens $\bm{i}^{c}_{t}$. The final input of the model is as follows:
$(\bm{a}_{t-1},({\color[rgb]{1,0,0}\bm{i}_{t}^{c}+\bm{z}^{c}_{t}},\bm{h}^{c}_{t},\bm{w}^{c}_{t})_{c=1}^{C},\bm{r}_{t},\bm{d}_{t})$ (3)
Unlike the standard encoder-decoder Transformer, our architecture does not perform cross-attention between the encoder and the decoder because information from the encoder has already been incorporated into the decoder through EMMA.
Model training.
We train the model to minimize cross-entropy loss with respect to the ground-truth (tokenized) trajectories in the training set. The label at each output position is the next token in the ground-truth sequence. In particular, we do not compute the losses at the positions of the action tokens and the first block’s tokens, because those tokens will be set during inference.
@@ 5.1 Baselines
We compare our model, which we call EMMA-LWM, with the followings:
@@
    Observational world model does not leverage textual information. It is identical to EMMA-LWM except that we zero out the manual representation $\bm{m}^{\textrm{enc}}$;
@@
    Standard is the encoder-decoder Transformer model following Vaswani et al. (2017) with multi-headed cross-attention between the decoder and the encoder. Similarly to EMMA-LWM, the model uses BERT to initially encode the manual into hidden vectors. The encoder applies self-attention to the hidden vectors of each description separately, instead of joining all vectors into a sequence and applying self-attention to it;
@@
    GPTHard is similar to EMMA-LWM but uses ChatGPT instead of EMMA to ground descriptions to entities. More details about this model are in Appendix A;
@@
    OracleParse is the same as GPTHard, but uses an oracle information extraction function. A description like “the crucial target is held by the wizard and the wizard is fleeing from you” is converted into “mage fleeing goal” for this model.
We train all models using AdamW (Loshchilov and Hutter, 2017) for $10^{5}$ iterations. For further details, please refer to Appendix C.
Table 1: Cross entropy losses ($\downarrow$) of different models on test ground-truth trajectories. Note that the minimum loss is non-zero because the Messenger environment is stochastic. We run each model with five different random seeds, selecting the final checkpoint for each seed based on the loss in the development NewAll split. We report the mean losses with 95% t-value confidence intervals. The bold number in each column indicates the best non-oracle mean.
@@ 5.2 Results
Table 1 shows the cross-entropy losses of all models on ground-truth trajectories sampled from the true environment dynamics (more in Appendix E). In the more difficult NewAttr and NewAll splits, our EMMA-LWM model consistently outperforms all baselines, nearing the performance of the OracleParse model. As expected, the Observational model is easily fooled by spurious correlations between identity and attributes, and among attributes. A specific example is illustrated in Figure 3. There, the Observational model incorrectly captures the movement of the whale and the queen. It also mistakenly portrays the whale as an enemy, whereas, in fact, the entity holds the message. In contrast, EMMA-LWM is capable of interpreting the previously unseen manual and accurately simulates the dynamics.
The performance of the Standard model is sensitive to initialization; in some runs, it performs as well as EMMA-LWM, but in others it performs as badly as Observational. A plausible explanation is that the model’s attention mechanism lacks sufficiently strong inductive biases to consistently find generalizable solutions. Our results agree with previous work on the lack of compositional generalizability of Transformers, which is often remedied by adding various forms of inductive bias (Keysers et al., 2020; Jiang and Bansal, 2021; Chaabouni et al., 2021; Dziri et al., 2023).
Another interesting finding is that the GPTHard model does not perform as well as expected. As a reminder, this model relies on ChatGPT to parse identities from descriptions and only needs to learn to extract attributes. Its underperformance compared to EMMA-LWM can be attributed to (i) the imperfection of ChatGPT in identifying identities in descriptions (its accuracy is around 90%; see Appendix B) and (ii) the fact that EMMA-LWM jointly learns to extract both identity and attribute words, which may be more effective than learning to extract only attribute words.
@@
In this evaluation, for each world model and test trajectory, we reset the model to the initial state of the trajectory and sequentially feed the actions in the trajectory to the model until it predicts the end of the episode. This process generates an imaginary trajectory. We refer to the evaluation trajectory as the real trajectory. We compute precisions of predicting non-zero rewards ($r\neq 0$) and terminations ($d=1$). To evaluate movement prediction, we compare the distances from the player to an entity in the real and imaginary trajectories. Concretely, let $\delta_{c,t}^{\text{real}}$ and $\delta_{c,t}^{\text{imag}}$ be the Hamming distances from the player to entity $c$ at the $t$-th time step in a real trajectory $\tau_{\textrm{real}}$ and an imaginary trajectory $\tau_{\textrm{imag}}$, respectively. We calculate the average difference in a specific time step: $\Delta_{\text{dist}}=\frac{1}{|\mathcal{D}_{\textrm{eval}}|}\sum_{\tau_{\textrm{real}}\in\mathcal{D}_{\textrm{eval}}}\frac{1}{T_{\textrm{min}}}\sum_{t=1}^{T_{\textrm{min}}}|\delta_{c,t}^{\text{real}}-\delta_{c,t}^{\text{imag}}|$ where $\mathcal{D}_{\textrm{eval}}$ is an evaluation split, $T_{\textrm{min}}=\min(|\tau_{\textrm{real}}|,|\tau_{\textrm{imag}}|)$, and $\tau_{\textrm{imag}}$ is generated from $\tau_{\textrm{real}}$ For example, for a chasing entity, $\delta_{c,t}^{\text{real}}$ decreases as $t$ increases. If a model mistakenly predicts the entity to be immobile, $\delta_{c,t}^{\text{imag}}$ remains a constant as $t$ progresses. In this case, $\Delta_{\text{dist}}$ is non-negligible, indicating an error. All evaluation metrics are given in Table 2. The ordering of the models is similar to that in the evaluation with ground-truth trajectories. EMMA-LWM is still superior to all baselines in all metrics.
@@ 5.3 Application: agents that discuss plans with humans
In this section, we showcase the practicality of our LWM by illustrating that it can facilitate plan discussions between an agent and a human supervisor. This approach has the potential to improve the transparency, safety, and performance of real-world agents.
We imagine an agent ordered to perform a task in a previously unseen environment (1(a)). Letting the agent perform the task immediately would be extremely risky because of its imperfect knowledge of the environment. Implementing a world model enables the agent to imagine a solution trajectory and present it to a human as a plan for review. Conveying plans as trajectories helps the human envision the future behavior of the agent in the real world. Furthermore, the human can improve this behavior by providing feedback to enhance the policy that produces the plan.
A human can update the policy by telling the agent which actions it should have taken. This type of feedback can be incorporated using some form of imitation learning. An agent equipped with a LWM additionally enables the human to update its policy by giving language feedback that aims to modify its world model. Although an observational world model also allows this form of adaptation, it requires much more effort from the human to generate the feedback. Concretely, the human has to generate observations in the same format as those in the agent’s plan (e.g., they have to draw grids in this setting). Furthermore, many abstract concepts may not be efficiently or precisely specified through non-verbal communication.
We simulate this scenario by placing agents with randomly initialized policies in test environments. These agents are forbidden to interact with the environments. However, they are equipped with world models, which allows for imaginary policy update. The world models are the ones we evaluated in the previous section. Importantly, the models were not trained on any data collected in the environments, simulating the fact that these environments are completely new to the agents.
We train all policies with imitation learning, considering two types of feedback: in online imitation learning (Ross et al., 2011), the expert suggests the best actions to take in the states present in the plan; in the filtered behavior cloning setting, the expert simply overwrites the agent’s plan with their own plan. In the latter setting, the agent chooses the plans that achieve the highest returns according to their world models to imitate. We experiment with a near-optimal expert and a suboptimal expert. We provide more details in Appendix D.
The agents endowed with LWMs can also process language feedback aiming to change their world models. This feedback is simulated by the game manuals accompanying the environments. It serves as the input $\bm{\ell}$ of the LWMs. We suppose that a human gives this feedback once to an agent, before adapting it via imitation learning.
We present the performance of the agents after adaptation in Table 3. Learning with the Observational world model amounts to the case where the human provides only imitation-learning feedback and cannot adapt the world model via language. Meanwhile, learning with EMMA-LWM represents the case where the human can use language feedback to improve the world model. In all evaluation settings, we observe significant improvements in the average return of policies that adopt our EMMA-LWM. There are still considerable gaps compared to using the OracleParse model, indicating that our model still has room for improvement.
Table 3: Average returns ($\uparrow$) in real environments of policies trained with imaginary imitation learning using world models. Bold numbers indicate the best non-oracle means in the corresponding settings. An expanded table with all models and details on how the metric was computed are available in Appendix E.